# PseudoNeg-MAE: Self-Supervised Point Cloud Learning using Conditional Pseudo-Negative Embeddings

### v1

@@ Abstract
We propose PseudoNeg-MAE, a novel self-supervised learning framework that enhances global feature representation of point cloud mask autoencoder by making them both discriminative and sensitive to transformations. Traditional contrastive learning methods focus on achieving invariance, which can lead to the loss of valuable transformation-related information. In contrast, PseudoNeg-MAE explicitly models the relationship between original and transformed data points using a parametric network COPE, which learns the localized displacements caused by transformations within the latent space. However, jointly training COPE with the MAE leads to undesirable trivial solutions where COPE outputs collapse to an identity. To address this, we introduce a novel loss function incorporating pseudo-negatives, which effectively penalizes these trivial invariant solutions and promotes transformation sensitivity in the embeddings. We validate PseudoNeg-MAE on shape classification and relative pose estimation tasks, where PseudoNeg-MAE achieves state-of-the-art performance on the ModelNet40 and ScanObjectNN datasets under challenging evaluation protocols and demonstrates superior accuracy in estimating relative poses. These results show the effectiveness of PseudoNeg-MAE in learning discriminative and transformation-sensitive representations.
I Introduction
@@
Self-Supervised Learning (SSL) has emerged as a crucial approach in point cloud analysis, allowing models to learn robust and generalizable representations from large volumes of unlabeled 3D data. With the growing accessibility of affordable 3D sensors and scanning devices, point clouds have become a widely used data format in applications such as autonomous driving, robotics, and augmented reality. SSL methods have achieved remarkable success in various downstream tasks involving point clouds, including 3D object classification [1, 2, 3, 4], 3D scene understanding [5, 6], and 3D part segmentation [1, 2, 3, 4]. Traditional SSL methods, particularly those based on contrastive learning, focus on learning invariant representations to input perturbations or transformations. However, this strict invariance to transformations can unintentionally ignore important information contained in the variations introduced by these transformations. Recent approaches [7, 8] address this limitation by introducing sensitivity to transformations, allowing models to explicitly encode the relationships between original and transformed point clouds in the latent space. Despite these advancements, the inherent design of their loss functions often leads to invariance as an inevitable outcome, making it challenging to learn representations highly sensitive to the transformations.
@@
In this work, we propose a novel self-supervised learning framework PseudoNeg-MAE, which effectively captures transformation-sensitive features while maintaining discriminative representations. Our method introduces a new loss function that leverages the conditional generation of pseudo negatives to prevent the model from collapsing to an invariant solution. By incorporating pseudo-negative embeddings corresponding to various transformations applied to the input point cloud, we explicitly model the localized displacements in the embedding space caused by different transformations. This approach ensures that the learned features remain sensitive to transformations while avoiding trivial solutions. To achieve this, we propose COnditional Pseudo Negatives Embedding network (COPE). COPE takes the transformation parameters as inputs and outputs a weight used to linearly project the original embedding. In this way, we conditionally generate pseudo-negatives depending on the set of random transformations. With pseudo-negatives, we regularize COPE through our novel loss function, ensuring that different transformations produce distinct embeddings in the feature space.
Our contrastive framework is built on the Mask AutoEncoder (MAE) transformer framework, combining the strengths of both paradigms. We employ a shared Siamese encoder with a transformer backbone, initialized with weights from an MAE. We train this architecture using our novel contrastive method, enhancing the global representations of the MAE to be more discriminative and sensitive to transformations. The transformer is required to generate global embeddings for both original and transformed inputs that satisfy the linear projection constraint imposed by COPE. Here the use of pseudo-negatives becomes crucial to avoid the trivial solution where COPE simply produces identity weights for all transformations.
@@
An advantage of our method is that it formulates relative pose estimation as an optimization problem in the embedding space, using transformation-sensitive features learned via the COPE. We introduce a novel inference algorithm that iteratively aligns source and target point cloud embeddings, achieving precise alignment even with large initial rotations. We validate the effectiveness of PseudoNeg-MAE through comprehensive experiments on shape classification and relative pose estimation tasks. On the ModelNet40 and ScanObjectNN datasets, our method achieves state-of-the-art performance, particularly under challenging protocols. For relative pose estimation tasks, our framework excels in accurately estimating relative rotations between point clouds, highlighting its capability to capture fine-grained transformation information.
II Related Works
@@
Self-Supervised Learning for Point Cloud: Self-Supervised Learning (SSL) has become increasingly prevalent for leveraging unlabeled data, especially in the challenging domain of 3D point clouds. The inherent disorder and lack of predefined structure in point clouds necessitate diverse strategies for extracting meaningful representations. Among these, contrastive methods [6, 9, 1, 10] exploit correspondences across different views of point clouds to establish unsupervised pretraining frameworks using InfoNCE loss [11, 12]. In parallel, recent advancements have introduced transformer-based masked autoencoder approaches, such as PointBERT [13], which decodes discrete tokens from point patches, and PointMAE [3] along with PointM2AE [4], which directly predict the masked point patches rather than tokens. Extending this paradigm, Point2Vec [2] employs the principles of Data2vec [14] to facilitate feature space reconstruction in point clouds. Diverging from earlier SSL techniques that mainly induce biases towards invariant feature learning or local relationships for reconstruction, our approach focuses on learning both discriminative and transformation-sensitive representations, significantly enhancing the performance across multiple downstream tasks.
@@
Sensitivity to Transformation in SSL methods: Contrastive learning  [12, 15, 16] has significantly advanced self-supervised representation learning; however, enforcing strict invariance to data augmentations can limit the expressive power of the learned features, especially in fine-grained tasks where details about specific transformations are crucial. To overcome this limitation, recent studies [17, 18, 7, 8] have explored enhancing models by incorporating sensitivity to transformations. One approach involves auxiliary tasks that require models to predict the transformations applied to the input data [19, 18, 17, 20], which helps preserve transformations-specific information within the representations and leads to less invariant feature spaces. However, these methods often lack guarantees of a consistent mapping between input transformations and changes in the latent representations [8]. Alternative approaches [21, 22], integrate loss functions that balance invariance with sensitivity to augmentations, enabling models to remain responsive to transformations while maintaining robustness. Additionally, recent methods [7, 8] have developed models that directly map transformations in the input space to transformations in the latent space, employing external predictors to modify representations accordingly. Despite these advancements, the inherent design of their loss functions often makes invariance an inevitable outcome [21, 8]. In our work, we address this issue by proposing a novel loss function that penalizes this trivial invariant feature spaces by leveraging the use of generated pseudo negatives. We demonstrate the effectiveness of incorporating an external predictor tailored to downstream tasks influenced by transformations. By introducing a specialized inference algorithm, we enhance model performance on tasks sensitive to these transformations. While previous methods largely focused on image-based classification, we extend these approaches to 3D point clouds, targeting both classification and relative pose estimation (i.e., registration).
Figure 1: Overview of the Proposed PseudoNeg-MAE Method: The method begins by processing a point cloud $\mathbf{x}_{i}$ and its transformed counterpart $\mathbf{x}_{i}^{+}$, extracting their global representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$, respectively, using a shared transformer architecture. Simultaneously, COPE network outputs the weight $\Theta_{g}$ and $\{\Theta_{g_{r}}\}$ based on the input and random transformations $g$ and $\{g_{r}\}$. $\{\Theta_{g_{r}}\}$ are used to generate pseudo-negative embeddings ${\{\tilde{\mathbf{z}}_{i}^{g_{r}}\}}$ which help in learning features that are both discriminative and sensitive to the transformation.
III Proposed Approach - PseudoNeg-MAE
@@
In this section, we introduce PseudoNeg-MAE, a method designed to learn feature representations that are both discriminative and sensitive to transformations. As shown in Fig. 1, we model the relationship between input transformations and displacements in the embedding space using linear projections. These projections are controlled by a network, COPE, which generates transformation-specific weights. We train the main network $f$ and COPE jointly. There is a risk that COPE could collapse into the identity matrix, producing invariant representations. To avoid this, we introduce a novel loss function that uses pseudo-negatives to regularize COPE and ensure that the projections are transformation-sensitive during training.
III-A Notations
@@
Let $f$ be a parametric network that maps an input space $\mathcal{X}$ to the unit sphere. Here $f$ is a composition of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, (i.e., $f=\mathcal{E}\circ\mathcal{D}$). Let $\mathcal{G}$ be a distribution of possible transformations in the input space $\mathcal{X}$ and $g$ denotes a transformation sampled from the distribution $\mathcal{G}$. For the given input point cloud $\mathbf{x}_{i}\in\mathcal{X}$, $g$ transforms it to a positive view $\mathbf{x}_{i}^{+}=g(\mathbf{x}_{i})$. The representations $\mathbf{z}$ of $\mathbf{x}_{i},\mathbf{x}_{i}^{+}$ is defined as $\mathbf{z}_{i}=f(\mathbf{x}_{i})\text{ and }\mathbf{z}_{i}^{+}=f(\mathbf{x}_{i}^{+})$. Both $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$ reside within the same representation space $\mathcal{Z}$ of $f$. We introduce an operator $u_{g}$ that models how the transformation $g$ affects the feature representation $\mathbf{z}_{i}$.
III-B Problem Statement
@@
Contrastive learning methods, such as InfoNCE [12], typically aim to learn a representation that is invariant to perturbations or transformations of the input data. This is achieved by encouraging similar instances to have similar representations in the embedding space, while dissimilar instances are pushed apart. This objective is equivalent to a combination of alignment $\mathcal{L}_{\text{align}}$ and uniformity $\mathcal{L}_{\text{unif}}$, as described in Eq. (3) below:
@@
$\mathcal{L}_{\text{align}}=$ $\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{z}_{i}^{+}\right\|_{2}^{2},$ (1)
$\mathcal{L}_{\text{ssl}}=$ $\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{unif}}.$ (3)
Specifically, $\mathcal{L}_{\text{align}}$ promotes the invariance property, while $\mathcal{L}_{\text{unif}}$ regularizes the learning process by preventing dimensional collapse, where $f(x)$ becomes constant for all $x$ [23]. Here $\tau$ is a temperature parameter.
@@
While invariant SSL models facilitate the learning process by discarding the variations introduced by transformations, they often neglect potentially critical information these variations might carry. To address this issue, a group of self-supervised learning methods [7, 8, 21] has introduced the concept of sensitivity to transformation, enabling the model to explicitly encode the relationships between the original and augmented data in latent space.
@@
These approaches introduce a modified alignment term, $\mathcal{L}_{\text{align}^{*}}=\frac{1}{N}\sum_{i=1}^{N}\left\|u_{g}(\mathbf{z}_{i})-\mathbf{z}_{i}^{+}\right\|_{2}^{2}$, where $u_{g}$ defines the relationship between the feature representations $\mathbf{z}_{i}$ and their augmented counterparts $\mathbf{z}i^{+}$ under the transformation $g$. When combined with the uniformity term from Eq. (2), this alignment term helps the model avoid dimensional collapse while enhancing its sensitivity to transformations.
However, a trivial solution to $\min_{f}\mathcal{L}_{\text{align}^{*}}$ is for all transformations $\forall g\in\mathcal{G}$ to collapse into the same embedding space (i.e., $u_{g}(\mathbf{z}_{i})=\mathbf{z}_{i}$), making the embeddings invariant to transformations. To avoid this, our method, PseudoNeg-MAE, introduces an additional constraint that produces localized displacements in the embedding space, ensuring that:
$u_{g_{r}}(\mathbf{z}_{i})\neq u_{g_{s}}(\mathbf{z}_{i});\quad$ $(g_{r},g_{s})\sim\mathcal{G}\times\mathcal{G}:g_{r}\neq g_{s}.$ (4)
III-C Sensitivity to Transformation through Pseudo Negatives
@@
To effectively model the relationship $u_{g}$, we introduce COnditional Pseudo Negatives Embedding (COPE) network based on hypernetworks [24, 8]. Specifically, COPE takes transformation parameter $g$ as an input and outputs the weight $\Theta_{g}\in\mathbb{R}^{d\times d}$. This weight is used as a linear projection on $\mathbf{z}_{i}$ which learns the localized displacements due to the transformation $g$.
$\Theta_{g}=\text{COPE}{}(g);\quad\mathbf{z}_{i}^{+}=\Theta_{g}\mathbf{z}_{i}.$ (5)
@@
A potential shortcut for COPE is to produce identity weights for all transformations, $\Theta_{g}=I$, effectively making the features invariant to transformations. This trivial solution undermines our goal of learning transformation-sensitive representations. To avoid this, we introduce the process of conditional generation of pseudo negatives using COPE. With pseudo negatives, we regularize COPE and prevent it from collapsing into the identity. This approach ensures that different transformations produce different projections, maintaining the sensitivity to the transformations. Specifically, we generate pseudo negatives by applying COPE to a set of randomly sampled transformations $\{g_{r}\}$ from $\mathcal{G}$, where $\{g_{r}\}$ are different from the input transformation $g$. For each $g_{r}$, we compute the corresponding weight $\Theta{g_{r}}$ with the COPE and generate the pseudo negative embedding $\mathbf{\tilde{{z}}}_{i}^{g_{r}}$:
$\mathbf{\tilde{{z}}}_{i}^{g_{r}}=\text{COPE}(g_{r})\mathbf{z}_{i}=\Theta_{g_{r}}\mathbf{z}_{i}.$ (6)
@@
$\mathcal{L}_{\text{cope}}=\log\left[\sum_{r=1}^{M}e^{-\left\|\Theta_{g}\mathbf{z}_{i}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+e^{-\left\|\Theta_{g}\mathbf{z}_{i}-\mathbf{z}_{i}^{+}\right\|_{2}^{2}/\tau}\right],$ (7)
where $M$ is the number of pseudo negatives. This loss penalizes the model for mapping different transformations to similar embeddings, thus preventing the trivial solution. Consider the case of perfect alignment where $\mathbf{z}_{i}^{+}=\Theta_{g}\mathbf{z}_{i}$. Minimizing $\mathcal{L}_{\text{cope}}$ is equivalent to optimizing $\log\left[\sum_{r=1}^{M}e^{-\left\|\Theta_{g}\mathbf{z}_{i}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+1\right]$, which effectively maximizes the pairwise distances between $\{\Theta_{g_{r}}\mathbf{z}_{i}\}$. Achieving this requires diverse $\{\Theta_{g_{r}}\}$ from the COPE, thereby penalizing trivial invariant solutions. We define the final loss $\mathcal{L}_{\text{PseudoNegMAE}}$ as follows:
@@
Specifically, $\mathcal{L}_{\text{unif}}$ encourages discriminative features and $\mathcal{L}_{\text{cope}}$ preserves the sensitivity to transformations within the feature space. This combination enables the model to learn discriminative and sensitive representations to input transformations. Fig. 2 illustrates the objective of our loss in terms of the anchor, the positive, and the pseudo-negatives. The loss $L_{\text{align}^{*}}$ encourages the positive to be pulled towards the anchor such that $\Theta_{g}$ satisfies the Eq. (5). $\mathcal{L}_{\text{cope}}$ encourages the pseudo negatives to be pushed away from the anchor. For a given $\mathbf{z}_{i}$ all the pseudo-negatives $\{\Theta_{g_{r}}\mathbf{z}_{i}\}$ only depend on the output of the COPE $\Theta_{g_{r}}$ network. This avoids $\Theta_{g}$ collapsing into identity, preventing the undesired invariant solution.
III-D COPE Network Architecture
COPE generates linear weights conditioned on the transformation set ${\{g}\}$. In our work, we consider rotations as transformations and utilize quaternions to represent them as inputs to the COPE. The network architecture is illustrated in the right section of Fig. 1.

### v2

@@ Abstract
We propose PseudoNeg-MAE, a novel self-supervised learning framework that enhances global feature representation of point cloud masked autoencoder by making them both discriminative and sensitive to transformations. Traditional contrastive learning methods focus on achieving invariance, discarding transformation-specific information. Recent approaches incorporate transformation sensitivity by explicitly modeling relationships between original and transformed inputs. However, they report an invariant-collapse phenomenon, where the predictor degenerates into identity mappings, resulting in latent representations that have limited variation across transformations. We propose a novel loss that explicitly penalizes invariant collapse, enabling the network to capture richer transformation cues while preserving discriminative representations. PseudoNeg-MAE uses a parametric network COPE, which learns the localized displacements caused by transformations within the latent space. However, jointly training COPE with the MAE leads to undesirable trivial solutions where COPE outputs collapse to an identity. To address this, we propose a loss that uses transformation-conditioned pseudo-negatives, to penalize such trivial invariant solutions. We validate PseudoNeg-MAE on shape classification and relative pose estimation tasks, where it achieves competitive performance on the ModelNet40 and ScanObjectNN datasets under challenging evaluation protocols and demonstrates superior accuracy in estimating relative poses compared to supervised methods.
11footnotetext: CSIRO Robotics, DATA61, CSIRO, Australia. E-mails: firstname.lastname@csiro.au22footnotetext: School of Electrical Engineering and Robotics, Queensland University of Technology (QUT), Brisbane, Australia.33footnotetext: Australian National University
Keywords: representation learning, point clouds, self-supervised learning
1 Introduction
@@
Self-Supervised Learning (SSL) has emerged as a crucial approach in point cloud analysis, allowing models to learn robust and generalizable representations from large volumes of unlabeled 3D data. With the growing accessibility of affordable 3D sensors and scanning devices, point clouds have become a widely used data format in applications such as autonomous driving, robotics, and augmented reality. SSL methods have achieved remarkable success in various downstream tasks involving point clouds, including 3D object classification [1, 2, 3, 4], 3D scene understanding [5, 6], and 3D part segmentation [1, 2, 3, 4]. Traditional SSL methods, particularly those based on contrastive learning, focus on learning invariant representations to input perturbations or transformations. However, enforcing strict invariance to transformations may suppress critical geometric cues, such as orientation or spatial displacement, that are essential for tasks in robotics where the structure and dynamics of the physical world must be preserved.
Recent approaches such as EquiMod [7] and SIE [8] address this limitation by introducing sensitivity to transformations (i.e., equivariant objective), allowing models to explicitly encode the relationships between original and transformed point clouds in the latent space. SIE [8] shows that the linear predictor used by EquiMod [7] can suppress transformation channels, collapse into an identity mapping, and produce representations that are no more informative than those of purely invariant methods. To mitigate this shortcut, SIE replaces the linear layer with a bias-free hypernetwork whose weights are conditioned on the transformation. Although the hypernetwork conditions its weights on the transformation, it cannot guarantee that the predicted transformation deviates meaningfully from the identity. This limitation constrains the model’s ability to capture significant local displacements in the latent space. Consequently, the resulting representations may exhibit only limited variation in response to different transformations, reducing their sensitivity. This underscores the need for strong regularization mechanisms that actively promote responsiveness to transformations.
@@
In this work, we propose a novel self-supervised learning framework, PseudoNeg-MAE, which learns transformation-sensitive representations while preserving discriminative capacity. Our key contribution is a loss function that bootstraps pseudo negatives conditioned on randomly sampled transformations, avoiding the risk of collapse to invariant solutions. These pseudo negatives serve as reference embeddings generated using the predictor, capturing the localized variations in the embedding space caused by different transformations. By incorporating them into the training objective, we encourage the model to produce representations that are sensitive to input transformations while avoiding degenerate solutions. To achive this, we introduce the COnditional Pseudo Negatives Embedding network (COPE). COPE takes transformation parameters as input and outputs a weight used to linearly project the original embeddings. This conditional mechanism enables the generation of transformation specific pseudo negatives. We use these pseudo negatives to regularize COPE through our proposed loss function, ensuring that distinct transformations yield sufficiently different embeddings in the latent space.
@@
An advantage of our method is that it formulates relative pose estimation as an optimization problem in the embedding space, using transformation-sensitive features learned via the COPE. We introduce a novel inference algorithm that iteratively aligns source and target point cloud embeddings, achieving precise alignment even with large initial misalignments. We validate the effectiveness of PseudoNeg-MAE through comprehensive experiments on shape classification and relative pose estimation tasks. On the ModelNet40 and ScanObjectNN datasets, our method achieves state-of-the-art performance, particularly under challenging protocols. For relative pose estimation tasks, our framework excels in accurately estimating relative poses between point clouds, leading to superior performance compared to supervised baselines.
2 Related Works
@@
Self-Supervised Learning for Point Cloud: Self-Supervised Learning (SSL) has become increasingly prevalent for leveraging unlabeled data, especially in the challenging domain of 3D point clouds. The inherent disorder and lack of predefined structure in point clouds necessitate diverse strategies for extracting meaningful representations. Among these, contrastive methods [6, 9, 1, 10] exploit correspondences across different views of point clouds to establish unsupervised pretraining frameworks using InfoNCE loss [11, 12]. In parallel, recent advancements have introduced transformer-based masked autoencoder approaches, such as PointBERT [13], which decodes discrete tokens from point patches, and PointMAE [3] along with PointM2AE [4], which directly predict the masked point patches rather than tokens. Extending this paradigm, Point2Vec [2] employs the principles of Data2vec [14] to facilitate feature space reconstruction in point clouds. Diverging from earlier SSL techniques that mainly induce biases towards invariant feature learning or local relationships for reconstruction, our approach focuses on learning both invariant and equivariant representations, significantly enhancing the performance across both downstream tasks.
Supervised Equivariant Networks: EPN [15] and E2PN [16] introduce SE(3)-equivariant architectures designed for point cloud analysis under supervised learning. EPN builds on KPConv [17] by incorporating group convolutions over discretized SO(3) to achieve rotation equivariance, while E2PN [16] improves efficiency by using quotient representations on $S^{2}\times\mathbb{R}^{3}$ with a permutation layer to recover full rotational structure. These models show strong performance on classification and relative pose estimation tasks but rely on labeled data. Additionally, Vector Neurons [18] represent another notable class of SO(3)-equivariant networks, using vector-valued features that transform linearly under rotation. EquivReg [19] leverages Vector Neurons for correspondence-free point cloud registration, solving for relative rotations in closed-form within the equivariant feature space. Our method instead leverages the properties of equivariance as a self-supervisory signal for representation learning with a non-equivariant backbone, allowing for flexible training without labeled data while still benefiting from transformation-aware learning. We compare against these supervised baselines to highlight the advantages of our self-supervised framework.
@@
Sensitivity to Transformation in SSL methods: Contrastive learning  [12, 20, 21] has significantly advanced self-supervised representation learning; however, enforcing strict invariance to data augmentations can limit the expressive power of the learned features, especially in robotic tasks where details about specific transformations are crucial. To overcome this limitation, recent studies [22, 23, 7, 8] have explored enhancing models by incorporating sensitivity to transformations. One approach involves auxiliary tasks that require models to predict the transformations applied to the input data [24, 23, 22, 25], which helps preserve transformations-specific information within the representations and leads to less invariant feature spaces. However, these methods often lack guarantees of a consistent mapping between input transformations and changes in the latent representations [8]. Alternative approaches [26, 27], integrate loss functions that balance invariance with sensitivity to augmentations, enabling models to remain responsive to transformations while maintaining robustness. Additionally, recent methods [7, 8] have developed models that directly map transformations in the input space to transformations in the latent space, employing external predictors to modify representations accordingly. These methods [7, 8] identify the invariant-collapse problem, where predictors degenerate into identity mappings, effectively removing essential transformation cues from the learned representations. SIE [8] addresses this by replacing the linear predictor with a transformation-conditioned hypernetwork. However, without strong regularization, the predicted transformations of SIE can still converge toward the identity, leading to limited transformation sensitivity. In contrast, we introduce a novel penalty loss based on transformation-conditioned pseudo-negatives that explicitly discourages collapse into invariant representations and promotes responsiveness to transformations. Since EquiMod [7] and SIE [8] were originally developed for the image domain, we include additional experiments in the supplementary material evaluating our method against these baselines on image-based tasks. The results confirm that our proposed penalty loss effectively mitigates invariant collapse and enhances transformation sensitivity.
3 Proposed Approach - PseudoNeg-MAE
@@
In this section, we introduce PseudoNeg-MAE, a method designed to learn feature representations that are both discriminative and sensitive to transformations. Our contrastive framework is built on the Masked AutoEncoder (MAE) transformer framework, combining the strengths of both predictive (MAE) and discriminative (contrastive) model training paradigms. As shown in Fig. 1, we model the relationship between input transformations and displacements in the embedding space using linear projections. These projections are controlled by a network, COPE, which generates transformation-specific weights. We train the main network $f$ and COPE jointly. There is a risk that COPE could collapse into the identity matrix, producing invariant representations. To avoid this, we introduce a novel loss function that uses pseudo-negatives to regularize COPE and ensure that the projections are equivariant during training.
Figure 1: Overview of PseudoNeg-MAE: It encodes a point cloud $\mathbf{x}_{i}$ and its transformed counterpart $\mathbf{x}_{i}^{+}$, extracting their global representations $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$, respectively, using a shared transformer architecture. Simultaneously, COPE network outputs the weight $\Theta_{g}$ and $\{\Theta_{g_{r}}\}$ based on the input and random transformations $g$ and $\{g_{r}\}$. $\{\Theta_{g_{r}}\}$ are used to generate pseudo-negatives ${\{\tilde{\mathbf{z}}_{i}^{g_{r}}\}}$.
@@
Let $f$ be a parametric network that maps an input space $\mathcal{X}$ to the unit sphere. Here $f$ is a composition of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, (i.e., $f=\mathcal{E}\circ\mathcal{D}$). Let $\mathcal{G}$ be a distribution of possible transformations in the input space $\mathcal{X}$, and $g$ denotes a transformation sampled from the distribution $\mathcal{G}$. For the given input point cloud $\mathbf{x}_{i}\in\mathcal{X}$, $g$ transforms it to a positive view $\mathbf{x}_{i}^{+}=g(\mathbf{x}_{i})$. The representations $\mathbf{z}$ of $\mathbf{x}_{i},\mathbf{x}_{i}^{+}$ is defined as $\mathbf{z}_{i}=f(\mathbf{x}_{i})\text{ and }\mathbf{z}_{i}^{+}=f(\mathbf{x}_{i}^{+})$. Both $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{+}$ reside within the same representation space $\mathcal{Z}$ of $f$. We introduce an operator $u_{g}$ that models how the transformation $g$ affects the feature representation $\mathbf{z}_{i}$.
@@
Contrastive learning methods, such as SimCLR [12], typically aim to learn a representation that is invariant to perturbations or transformations of the input data. This is achieved by encouraging similar instances to have similar representations in the embedding space, while dissimilar instances are pushed apart. This objective is equivalent to a combination of alignment $\mathcal{L}_{\text{align}}$ and uniformity $\mathcal{L}_{\text{unif}}$, as described in Eq. (1) below:
@@
$\mathcal{L}_{\text{align}}=$ $\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{z}_{i}^{+}\right\|_{2}^{2};\mathcal{L}_{\text{unif}}=\log\mathbb{E}_{(\mathbf{z}_{i},\mathbf{z}_{k})\sim\mathcal{Z}\times\mathcal{Z}}e^{-\left\|\mathbf{z}_{i}-\mathbf{z}_{k}\right\|_{2}^{2}/\tau};\mathcal{L}_{\text{ssl}}$ $=\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{unif}}.$ (1)
@@
Specifically, $\mathcal{L}_{\text{align}}$ promotes the invariance property, while $\mathcal{L}_{\text{unif}}$ regularizes the learning process by preventing dimensional collapse, where $f(x)$ becomes constant for all $x$ [28]. Here $\tau$ is a temperature parameter. While invariant SSL models facilitate the learning process by discarding the variations introduced by transformations, they often neglect potentially critical information that these variations might carry. To address this issue, a group of self-supervised learning methods [7, 8, 26] has introduced the concept of sensitivity to transformation, enabling the model to explicitly encode the relationships between the original and augmented data in the latent space.
@@
These approaches introduce a modified alignment term, $\mathcal{L}_{\text{align}^{*}}=\frac{1}{N}\sum_{i=1}^{N}\left\|u_{g}(\mathbf{z}_{i})-\mathbf{z}_{i}^{+}\right\|_{2}^{2}$, where $u_{g}$ defines the relationship between the feature representations $\mathbf{z}_{i}$ and their augmented counterparts $\mathbf{z}_{i}^{+}$ under the transformation $g$. When combined with the uniformity term from Eq. (1), this alignment term helps the model avoid dimensional collapse while enhancing its sensitivity to transformations. A trivial solution to $\min_{f}\mathcal{L}_{\text{align}^{*}}$ is for all transformations $\forall g\in\mathcal{G}$ to collapse into the same embedding space (i.e., $u_{g}(\mathbf{z}_{i})=\mathbf{z}_{i}$), making the embeddings invariant to transformations. To avoid this, our method, PseudoNeg-MAE, introduces an additional constraint that produces localized displacements in the embedding space, ensuring that: $u_{g_{r}}(\mathbf{z}_{i})\neq u_{g_{s}}(\mathbf{z}_{i});\quad(g_{r},g_{s})\sim\mathcal{G}\times\mathcal{G}:g_{r}\neq g_{s}.$
@@
Sensitivity to Transformation through Pseudo Negatives: To effectively model the relationship $u_{g}$, we introduce COnditional Pseudo Negatives Embedding (COPE) network based on hypernetworks [29, 8]. Specifically, COPE takes transformation parameter $g$ as an input and outputs the weight $\Theta_{g}=\text{COPE}(g)\in\mathbb{R}^{d\times d}$. This weight is used as a linear projection on $\mathbf{z}_{i}$ which learns the localized displacements due to the transformation $g$.
@@
A potential shortcut for COPE is to produce identity weights for all transformations, $\Theta_{g}=I$, making the features invariant to transformations. This trivial solution undermines our goal of learning transformation-sensitive representations. To avoid this, we introduce the process of conditional generation of pseudo negatives using COPE. With pseudo negatives, we regularize COPE and prevent it from collapsing into the identity. This approach ensures that different transformations produce different projections in the embedding, maintaining the sensitivity to the transformations. Specifically, we generate pseudo negatives by applying COPE to a set of randomly sampled transformations $\{g_{r}\}$ from $\mathcal{G}$, where $\{g_{r}\}$ are different from the input transformation $g$. For each $g_{r}$, the corresponding weight $\Theta{g_{r}}$ with the COPE is computed and generated the pseudo negative embedding $\mathbf{\tilde{{z}}}_{i}^{g_{r}}$:
$\mathbf{\tilde{z}}_{i}^{g_{r}}=\frac{\text{COPE}(g_{r})\mathbf{z}_{i}}{\|\text{COPE}(g_{r})\mathbf{z}_{i}\|}=\frac{\Theta_{g_{r}}\mathbf{z}_{i}}{\|\Theta_{g_{r}}\mathbf{z}_{i}\|}.$ (2)
@@
$\mathcal{L}_{\text{cope}}=\frac{1}{N}\sum_{i=1}^{N}\log\left[\sum_{r=1}^{M}e^{-\left\|\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+e^{-\left\|\frac{\Theta_{g}\mathbf{z}_{i}}{{\|\Theta_{g}\mathbf{z}_{i}\|}}-\mathbf{z}_{i}^{+}\right\|_{2}^{2}/\tau}\right],$ (3)
where $M$ is the number of pseudo negatives. This loss penalizes the model for mapping different transformations to similar embeddings, thus preventing the trivial solution. Consider the case of perfect alignment where $\mathbf{z}_{i}^{+}=\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}$. Minimizing $\mathcal{L}_{\text{cope}}$ is equivalent to optimizing $\log\left[\sum_{r=1}^{M}e^{-\left\|\frac{\Theta_{g}\mathbf{z}_{i}}{{\|\Theta_{g}\mathbf{z}_{i}\|}}-\mathbf{\tilde{{z}}}_{i}^{g_{r}}\right\|_{2}^{2}/\tau}+1\right]$, which effectively maximizes the pairwise distances between $\{\Theta_{g_{r}}\mathbf{z}_{i}\}$. Achieving this requires diverse $\{\Theta_{g_{r}}\}$ from the COPE, thereby penalizing trivial invariant solutions. We define the final loss $\mathcal{L}_{\text{PseudoNegMAE}}$ as follows:
@@
Specifically, $\mathcal{L}_{\text{unif}}$ encourages discriminative features and $\mathcal{L}_{\text{cope}}$ preserves the sensitivity to transformations within the feature space. This combination enables the model to learn discriminative and sensitive representations to input transformations. The loss $L_{\text{align}^{*}}$ encourages the positive $\mathbf{z}_{i}$ to be pulled towards the anchor. $\mathcal{L}_{\text{cope}}$ encourages the pseudo negatives $\mathbf{\tilde{{z}}}_{i}^{g_{r}}$ to be pushed away from the anchor $\frac{\Theta_{g}\mathbf{z}_{i}}{\|\Theta_{g}\mathbf{z}_{i}\|}$. For a given $\mathbf{z}_{i}$ all the pseudo-negatives $\{\Theta_{g_{r}}\mathbf{z}_{i}\}$ only depend on the output of the COPE $\Theta_{g_{r}}$ network. This avoids $\Theta_{g}$ collapsing into identity, preventing the undesired invariant solution. Unlike SIE [8] whose equivariance is mainly enforced through an alignment loss, a term that trivially vanishes when the predictor collapses to the identity, $\mathcal{L}_{COPE}$ conditioned with the Pseudo-Negatives reaches its maximum when predictor collapse to identity. The full proof appears in the Supplementary Materials.